# Smallness of neutrino masses and leptogenesis in

### v1

@@ Abstract
We consider 331 composite Higgs model (CHM3) in which the Lagrangian of the strongly coupled sector is invariant with respect to global $SU(3)_{C}\times SU(3)\times U(1)_{6}$ symmetry that can originate from $SU(6)$ subgroup of $E_{6}$ and contains the gauge group of the standard model (SM) as a subgroup. The breakdown of the approximate $SU(3)\times U(1)_{6}$ symmetry down to $SU(2)_{W}\times U(1)_{Y}$ subgroup around the scale $f\sim 10\,\mbox{TeV}$ results in a set of pseudo–Nambu–Goldstone bosons (pNGBs) that, in particular, involves Higgs doublet. The generation of the masses of the SM fermions in the CHM3 is discussed. We argue that an approximate discrete $Z_{2}$ symmetry may give rise to tiny masses of the left–handed neutrinos and several composite fermions with masses $1-2\,\mbox{TeV}$. The lepton and baryon asymmetries can be generated within the CHM3 via the out–of equilibrium decays of extra Majorana particle into the Higgs doublet and these composite fermions.
@@ 1 Introduction
The smallness of three neutrino masses may be naturally explained within seesaw models [1]. In these models the lepton asymmetry can be induced via the out–of equilibrium decays of the lightest right–handed neutrino [2] since all three Sakharov conditions [3] are satisfied in this case. This asymmetry gets partially converted into baryon asymmetry via sphaleron processes [4].
@@
The Lagrangian of the strongly interacting sector of the minimal composite Higgs model (MCHM) [8, 9] possesses an approximate global $SU(3)_{C}\times SO(5)\times U(1)_{X}$ symmetry. $SU(3)_{C}\times SU(2)_{W}\times U(1)_{Y}$ gauge group of the SM is a subgroup of $SU(3)_{C}\times SO(5)\times U(1)_{X}$. Near the scale $f$ the approximate $SO(5)\times U(1)_{X}$ symmetry is expected to be broken down to $SO(4)\times U(1)^{\prime}_{X}\cong SU(2)_{W}\times SU(2)_{R}\times U(1)^{\prime}_{X}$. Such breakdown results in four pseudo–Nambu–Goldstone bosons (pNGBs) that compose the Higgs doublet $H$. The custodial symmetry $SU(2)_{cust}\subset SO(4)\cong SU(2)_{W}\times SU(2)_{R}$ [10] protects the Peskin–Takeuchi $\hat{T}$ parameter [11] against new physics contributions. The contributions of new bound states to the EW observables in the CHMs were analysed in Refs. [12][20]. The implications of these models were also examined for Higgs physics [15][16], [21][24], gauge coupling unification [25][26], dark matter [13], [22], [26][27] and collider phenomenology [14][15], [17], [24], [28][30]. Different aspects of non-minimal CHMs were explored in Refs. [13], [21][22], [26][27], [31].
The SM bosons and fermions except the Higgs doublet are superpositions of elementary states and their composite partners. Within partial compositeness framework [32, 33] the couplings of the SM states to the Higgs doublet are determined by the fractions of their compositeness. The mass hierarchy in the quark and charged lepton sectors can be reproduced in this case if the fractions of compositeness of all fermions except the top quark are rather small. This also leads to the partial suppression of flavour–changing processes [32]. The non–diagonal flavour transitions in the quark and lepton sectors within the CHMs were examined in Refs. [18][20], [28], [34][35] and [29], [35][37], respectively. It was shown that in general the compositeness scale $f$ in these models is required to be larger than $10\,\mbox{TeV}$ [18][19], [28], [34], [36]. However in the CHMs with extra flavour symmetries the phenomenologically acceptable scenarios can be obtained even for $f\sim 1\,\mbox{TeV}$ [17][18], [28][29], [35], [38].
@@
The paper is organised as follows. In the next section we specify the CHM3, discuss the generation of masses of the SM fermions and argue that an approximate discrete $Z_{2}$ symmetry can lead to tiny masses of the left–handed neutrino. This discrete symmetry may also result in relatively light composite partners of the left–handed leptons which might be detected at the LHC in the near future. The leptogenesis in such scenario is considered in section 3. Our results are summarised in section 4.
2 CHM3 with approximate $Z_{2}$ symmetry
@@
When $f$ is much larger than the vacuum expectation value of the Higgs field $v\simeq 246\,\mbox{GeV}$ the first two components of $\Omega$ can be identified with the SM–like Higgs doublet while the third component involves the SM singlet scalar $\phi_{0}$. In the CHM3 the $U(1)_{Y}$ weak hypercharges $Y_{i}$ are linear combinations
@@
In order to ensure that the left–handed neutrino states gain masses which are much smaller than the charged lepton masses an approximate $Z_{2}$ discrete symmetry is imposed. All multiplets except ${\bf\overline{3}}^{i}_{-1/3}$ are even under $Z_{2}$ while ${\bf\overline{3}}^{i}_{-1/3}$ are $Z_{2}$ odd. Therefore the operators ${\bf\overline{3}}^{i}_{+2/3}\Omega^{\dagger}{\bf\overline{3}}^{i}_{-1/3}$ are strongly suppressed and only interactions
@@
On the other hand the interactions $(L_{1i}H)(L_{1j}H)$, that could induce the masses of the left–handed neutrino, are forbidden by the $U(1)_{6}$ symmetry. The appropriate interactions in the strongly coupled sector
@@
which are not suppressed by neither $U(1)_{6}$ nor $Z_{2}$ symmetry. The $SU(6)$ structure, that leads to the interactions (12)-(15), can be written as
@@
where $c_{11i}\approx c_{22i}\approx 1$ while $s_{1i}$, $s_{2i}$, $s_{12i}$ and $s_{21i}$ are quite small. The observed mass hierarchy in the lepton sector implies that $s_{2i}\ll s_{1i}$. The smallness of $s_{2i}$ can be caused by the approximate discrete $Z_{2}$ symmetry. In this case $\tilde{L}_{2i}$ can be also much lighter than $f$. The scenarios with relatively light $\tilde{L}_{2i}$ are going to be considered in the next section.
The approximate $Z_{2}$ symmetry may also lead to the observed mass hierarchy in the lepton sector within the MCHM. Nevertheless from Eq. (5) it follows that the mass terms of the components of ${\bf 5}^{i}_{0}$ are not suppressed by the approximate $Z_{2}$ symmetry. As a consequence they can gain masses of the order of $f$.
@@ 3 Leptogenesis in the CHM3
Hereafter we explore the CHM3 with approximate $Z_{2}$ symmetry that leads to three generations of relatively light composite resonances $\tilde{L}_{2i}$. The SM singlet components $N_{i}$ of the corresponding $SU(3)$ multiplets ${\bf\overline{3}}^{i}_{-1/3}$ gain masses around $f$ through the interactions (15) while the $SU(2)_{W}$ doublet components of these multiplets are substantially lighter than $f$. This scenario is realized in Nature if the compositeness scale $f\gtrsim 10\,\mbox{TeV}$. In this case the masses of $\tilde{L}_{2i}$ may vary between $1-2\,\mbox{TeV}$ so that it might be possible to discover such fermion states in the near future at the LHC. The couplings of these composite resonances to the SM leptons are suppressed because they are determined by the small parameters $s_{2i}$.
Now suppose that the weakly–coupled sector includes an additional elementary Majorana fermion $n$ which mixes with the SM singlet resonances $N_{i}$. The corresponding mixing parameters are expected to be of order of $s_{2i}\lesssim 10^{-5}$. If $n$ and $N_{i}$ have masses around $10\,\mbox{TeV}$ they can decay into leptons and Higgs doublet inducing lepton asymmetry. We further assume that $N_{i}$ are somewhat heavier than the elementary fermion $n$. The part of the CHM3 Lagrangian, which describes the interactions of $n=N_{0}$ and $N_{i}$ with the SM leptons, $\tilde{L}_{2i}$ and Higgs doublet $H$, can be presented as a sum
@@
where $i,j=1,2,3$ and $x=0,1,2,3$. Since $\ell_{i}$ and $N_{0}$ are mainly elementary fields $|g_{i0}|\ll|g_{ij}|$, $|g_{ij}|\sim|h_{j0}|\ll|h_{ij}|$. At the same time the Yukawa couplings $h_{ij}$ are not suppressed by the $Z_{2}$ symmetry and therefore $|h_{ij}|\gtrsim 0.1$.
@@
It is convenient to introduce $Y_{\Delta B}$ which is the baryon asymmetry relative to the entropy density. The observed value of $Y_{\Delta B}$ is given by
@@
Using an approximate formula [42]
$Y_{\Delta B}\sim 10^{-3}\,\varepsilon_{0}\,\eta_{0}\,$, (23)
one can estimate the induced baryon asymmetry. In Eq. (23) $\eta_{0}$ is an efficiency factor that varies from 0 to 1. Without washout effects the decays of the Majorana fermion $n$ would result in $\eta_{0}=1$. The washout processes reduce the generated baryon asymmetry by the factor $\eta_{0}$. Here we concentrate on the so–called strong washout scenario when the efficiency factor $\eta_{0}$ can be estimated as
$\eta_{0}\simeq H(T=M_{1})/\Gamma\,$$\qquad\qquad\qquad$
@@
where $H$ is the Hubble expansion rate and $g_{*}=n_{b}+\dfrac{7}{8}\,n_{f}$ is the number of relativistic degrees of freedom in the thermal bath. In the scenario under consideration $g_{*}=128.75$.
From Eqs. (24) it follows that $\eta_{0}$ diminishes with increasing $|h_{10}|$. The corresponding dependence is shown in Fig. 1a. In Eq. (23) both the decay asymmetry $\varepsilon_{0}$ and the efficiency factor are smaller than unity. Thus the appropriate baryon asymmetry can be generated only if $\eta_{0}\gtrsim 10^{-7}$. The results of the numerical analysis presented in Fig. 1a demonstrate that for $M_{n}\simeq 10\,\mbox{TeV}$ this condition can be fulfilled when $|h_{10}|\lesssim 0.001$. Otherwise the absolute value of $Y_{\Delta B}$ tends to be too small. Within the CHM3 the value of $|h_{10}|$ is expected to be of the order of $s_{2i}$ because fermion $n$ is mainly elementary state, i.e. $|h_{10}|\lesssim 10^{-5}$.
(a)                        (b)
Figure 1: Logarithm (base 10) of the absolute values of the efficiency factor $\eta_{0}$ and decay asymmetry $\varepsilon_{0}$ for $h_{20}=h_{30}=0$. In (a) the absolute value of $\eta_{0}$ is given as a function of logarithm (base 10) of $|h_{10}|$ for $M_{n}=5\,\mbox{TeV}$ (dashed–dotted line), $M_{n}=10\,\mbox{TeV}$ (solid line) and $M_{n}=15\,\mbox{TeV}$ (dashed line). In (b) the absolute value of $\varepsilon_{0}$ is presented as a function of logarithm (base 10) of $\Delta\varphi_{1}=\Delta\varphi_{2}=\Delta\varphi_{3}=\Delta\varphi$ for $M_{n}=10\,\mbox{TeV}$, $M_{1}=12\,\mbox{TeV}$, $M_{2}=14\,\mbox{TeV}$, $M_{3}=16\,\mbox{TeV}$, $|h_{11}|=|h_{12}|=|h_{13}|=|h|$, $|h|=0.1$ (solid line) and $|h|=0.3$ (dashed line).
Since in the scenario under consideration the efficiency factor is not negligibly small, i.e. $\eta_{0}\sim 0.01$, the absolute value of $Y_{\Delta B}$ is determined by the decay asymmetry $\varepsilon_{0}$. From Eq. (21) one can see that $\varepsilon_{0}$ is set by $|h_{1j}|$ and CP violating phases $\Delta\varphi_{j}$. It also depends on the ratio of masses of Majorana fermions $M_{j}/M_{n}$. Because the Yukawa couplings of $N_{j}$ to $L_{2i}$ and the Higgs doublet are not suppressed by the approximate $Z_{2}$ symmetry, all $|h_{ij}|$ should be relatively large, i.e. $|h_{ij}|\gtrsim 0.1$. Here we fix $M_{n}=10\,\mbox{TeV}$, $M_{1}=12\,\mbox{TeV}$, $M_{2}=14\,\mbox{TeV}$, $M_{3}=16\,\mbox{TeV}$, $|h_{11}|=|h_{12}|=|h_{13}|=h$ and $\Delta\varphi_{1}=\Delta\varphi_{2}=\Delta\varphi_{3}=\Delta\varphi$. In Fig. 1b the dependence of $|\varepsilon_{0}|$ on $\Delta\varphi$ for different values of $h$ is shown. The absolute value of the decay asymmetry grows monotonically with increasing of $|h|$. The results presented in Fig. 1b indicate that for $h\gtrsim 0.1$ and $\eta_{0}\sim 0.01$ the observed baryon asymmetry can be obtained for small CP violating phases, i.e. $\Delta\varphi\ll 0.01$.

### v2

@@ Abstract
We consider 331 composite Higgs model (CHM3) in which the Lagrangian of the strongly coupled sector is invariant with respect to global $SU(3)_{C}\times SU(3)\times U(1)_{6}$ symmetry that can originate from $SU(6)$ subgroup of $E_{6}$ and contains the gauge group of the standard model (SM) as a subgroup. The breakdown of the approximate $SU(3)\times U(1)_{6}$ symmetry down to $SU(2)_{W}\times U(1)_{Y}$ subgroup around the scale $f\sim 10\,\mbox{TeV}$ results in a set of pseudo–Nambu–Goldstone bosons (pNGBs) that, in particular, involves Higgs doublet. The generation of the masses of the SM fermions in the CHM3 is discussed. We argue that approximate $U(1)_{L}$ and discrete $Z_{2}$ symmetries may give rise to tiny masses of the left–handed neutrinos and several composite fermions with masses $1-2\,\mbox{TeV}$. The lepton and baryon asymmetries can be generated within the CHM3 via the out–of equilibrium decays of extra Majorana particle into the Higgs doublet and these composite fermions.
@@ 1 Introduction
The smallness of three neutrino masses may be naturally explained within seesaw models [1]. In these models the lepton asymmetry can be induced via the out–of equilibrium decays of the lightest right–handed neutrino [2] since all three Sakharov conditions [3] are satisfied in this case. This asymmetry gets partially converted into a baryon asymmetry via sphaleron processes [4].
@@
The Lagrangian of the strongly interacting sector of the minimal composite Higgs model (MCHM) [8, 9] possesses an approximate global $SU(3)_{C}\times SO(5)\times U(1)_{X}$ symmetry. $SU(3)_{C}\times SU(2)_{W}\times U(1)_{Y}$ gauge group of the SM is a subgroup of $SU(3)_{C}\times SO(5)\times U(1)_{X}$. Near the scale $f$ the approximate $SO(5)\times U(1)_{X}$ symmetry is expected to be broken down to $SO(4)\times U(1)_{X}\cong SU(2)_{W}\times SU(2)_{R}\times U(1)_{X}$. Such breakdown results in four pseudo–Nambu–Goldstone bosons (pNGBs) that compose the Higgs doublet $H$. The custodial symmetry $SU(2)_{cust}\subset SO(4)\cong SU(2)_{W}\times SU(2)_{R}$ [10] protects the Peskin–Takeuchi $\hat{T}$ parameter [11] against new physics contributions. The contributions of new bound states to the EW observables in the CHMs were analysed in Refs. [12][20]. The implications of these models were also examined for Higgs physics [15][16], [21][24], gauge coupling unification [25][26], dark matter [13], [22], [26][27] and collider phenomenology [14][15], [17], [24], [28][30]. Different aspects of non-minimal CHMs were explored in Refs. [13], [21][22], [26][27], [31].
The SM bosons and fermions except the Higgs doublet are superpositions of elementary states and their composite partners. Within the partial compositeness framework [32, 33] the couplings of the SM states to the Higgs doublet are determined by the fractions of their compositeness. The mass hierarchy in the quark and charged lepton sectors can be reproduced in this case if the fractions of compositeness of all fermions except the top quark are rather small. This also leads to the partial suppression of flavour–changing processes [32]. The non–diagonal flavour transitions in the quark and lepton sectors within the CHMs were examined in Refs. [18][20], [28], [34][35] and [29], [35][37], respectively. It was shown that in general the compositeness scale $f$ in these models is required to be larger than $10\,\mbox{TeV}$ [18][19], [28], [34], [36]. However in the CHMs with extra flavour symmetries the phenomenologically acceptable scenarios can be obtained even for $f\sim 1\,\mbox{TeV}$ [17][18], [28][29], [35], [38].
@@
The paper is organised as follows. In the next section we specify the CHM3, discuss the generation of masses of the SM fermions and argue that approximate $U(1)_{L}$ and discrete $Z_{2}$ symmetries can lead to tiny masses of the left–handed neutrino. The approximate $Z_{2}$ symmetry may also result in relatively light composite partners of the left–handed leptons which might be detected at the LHC in the near future. The leptogenesis in such scenario is considered in section 3. Our results are summarised in section 4.
2 CHM3 with approximate $U(1)_{L}$ and $Z_{2}$ symmetries
@@
When $f$ is much larger than the vacuum expectation value of the Higgs field $v\simeq 246\,\mbox{GeV}$ the first two components of $\Omega$ can be identified with the SM–like Higgs doublet while the third component involves the SM singlet scalar $\phi_{0}$. The masses of all composite resonances except the Higgs doublet in the model under consideration are set by the scale $f\gtrsim 5-6\,\mbox{TeV}$. For instance, the SM singlet scalar $\phi_{0}$ tends to acquire a mass in the multi TeV range. Thus all new states predicted by the CHM3 are rather heavy and decouple from the SM spectrum.
In the CHM3 the $U(1)_{Y}$ weak hypercharges $Y_{i}$ are linear combinations
@@
Multiplets ${\bf\overline{3}}^{i}_{-1/3}$ and ${\bf 3}^{i}_{+1/3}$ get combined forming vectorlike fermion states. In order to ensure that the left–handed neutrino states gain masses, which are much smaller than the charged lepton masses, approximate $U(1)_{L}$ and discrete $Z_{2}$ symmetries are imposed. All multiplets except ${\bf 3}^{i}_{+1/3}$ are even under approximate $Z_{2}$ while ${\bf 3}^{i}_{+1/3}$ are $Z_{2}$ odd. The $U(1)_{L}$ symmetry is associated with the lepton number conservation. Multiplets ${\bf\overline{3}}^{i}_{+2/3}$ and ${\bf 3}^{i}_{-2/3}$ carry $U(1)_{L}$ charges (1)(-1) and $(+1)$ respectively whereas all other multiplets of the composite states have zero $U(1)_{L}$ charges. The Lagrangian of the strongly coupled sector respects the $U(1)_{L}$ global symmetry. Therefore the operators ${\bf\overline{3}}^{i}_{+2/3}\Omega^{\dagger}{\bf\overline{3}}^{i}_{-1/3}$ are forbidden and only interactions
@@
On the other hand the interactions $(L_{1i}H)(L_{1j}H)$ and $(L_{1i}H)(L_{2j}H)$, that could induce the masses of the left–handed neutrino, are forbidden by the $U(1)_{L}$ symmetry. The appropriate interactions in the strongly coupled sector
@@
which are not suppressed by neither $U(1)_{L}$ nor $Z_{2}$ symmetry. The $SU(6)$ structure, that leads to the interactions (12) and (15), can be written as
@@
where $c_{11i}\approx c_{22i}\approx 1$ while $s_{1i}$, $s_{2i}$, $s_{12i}$ and $s_{21i}$ are quite small. Substituting $L_{2i}$ from Eqs. (17) into Eq. (14) one obtains the following leading order analytic expression for the neutrino mass matrix ($\mathcal{M}^{\nu}_{ij}$)
$\mathcal{M}^{\nu}_{ij}=-\dfrac{\varkappa_{ij}}{2f}s_{2i}s_{2j}v^{2}\,.$
From Eqs. (17) and (13) it follows that the masses of the charged leptons are determined by $Y^{e}_{ij}\sim 1$, $s_{1i}$ and the fractions of the compositeness of the charged right–handed leptons $s_{ei}$. All measured charged lepton masses can be reproduced when $|s_{1i}|$ and $|s_{ei}|$ vary between $0.001$ and $0.1$ . Since $\varkappa_{ij}\sim 1$, the appropriate neutrino mass scale ($\sim 0.1\,\mbox{eV}$) requires that $s_{2i}\lesssim 10^{-5}$, i.e. $s_{2i}\ll s_{1i}$.
The smallness of $s_{2i}$ can be caused by both $U(1)_{L}$ and approximate $Z_{2}$ symmetries. If the mixing between $\tilde{L}_{2i}$ and elementary states vanishes then $s_{12i}=s_{21i}=s_{2i}=0$ and the Lagrangian of the CHM3 possesses an exact global $U(1)_{L}$ symmetry because in the strongly interacting sector the lepton number violating operators cannot be induced. When $Z_{2}$ symmetry is exact, $\tilde{L}_{2i}$ remain massless and the values of $s_{12i}$, $s_{21i}$ and $s_{2i}$ go to zero. The approximate $Z_{2}$ symmetry implies that $s_{2i}$ are small and $\tilde{L}_{2i}$ are much lighter than $f$. The scenarios with relatively light $\tilde{L}_{2i}$ are going to be considered in the next section.
The approximate $Z_{2}$ symmetry may also lead to rather small values of $s_{2i}$ in the MCHM. Nevertheless from Eq. (5) it follows that the mass terms of the components of ${\bf 5}^{i}_{0}$ are not suppressed by the approximate $Z_{2}$ symmetry. As a consequence they can gain masses of the order of $f$.
@@ 3 Leptogenesis in the CHM3
Hereafter we explore the CHM3 with an approximate $Z_{2}$ symmetry that leads to three generations of relatively light composite resonances $\tilde{L}_{2i}$. The SM singlet components $N_{i}$ of the corresponding $SU(3)$ multiplets ${\bf\overline{3}}^{i}_{-1/3}$ gain masses around $f$ through the interactions (15) while the $SU(2)_{W}$ doublet components of these multiplets are substantially lighter than $f$. This scenario is realized in Nature if the compositeness scale $f\gtrsim 10\,\mbox{TeV}$. In this case the masses of $\tilde{L}_{2i}$ may vary between $1-2\,\mbox{TeV}$ so that it might be possible to discover such fermion states in the near future at the LHC. The couplings of these composite resonances to the SM leptons are suppressed because they are determined by the small parameters $s_{2i}$.
As mentioned before, if $f\gtrsim 10\,\mbox{TeV}$ one does not need to impose any extra flavour symmetries to suppress the non–diagonal flavour transitions in the CHMs. Here we assume that the low energy effective Lagrangian of the CHM3 is invariant with respect to a $Z^{B}_{2}$ symmetry, which is a discrete subgroup of $U(1)_{B}$, i.e.
$Z^{B}_{2}=(-1)^{3B}\,$,
where $B$ is the baryon number. This symmetry provides a mechanism for the suppression of the operators which lead to rapid proton decay. Other processes that violate baryon number, most notably neutron–antineutron transitions, are sufficiently strongly suppressed by the relatively large value of $f\gtrsim 10\,\mbox{TeV}$ as well as the rather small mixing between elementary states and their composite partners. Meanwhile for so large $f$ a substantial degree of tuning is required to get a $125\,\mbox{GeV}$ Higgs state. Consequently, here we focus on the scenarios with $f\simeq 10\,\mbox{TeV}$.
Now suppose that the weakly–coupled sector includes an additional elementary Majorana fermion which mixes with the SM singlet components of ${\bf\overline{3}}^{i}_{-1/3}$. In general, since such mixing does not violate neither $U(1)_{L}$ nor $Z_{2}$ symmetry, the compositeness fractions of the corresponding Majorana mass eigenstate $n$ and other SM fermions are expected to be of the same order of magnitude, i.e. $\sim|s_{1i}|\sim|s_{ei}|$. If $n$ and $N_{i}$ have masses around $10\,\mbox{TeV}$ they can decay into leptons and Higgs doublet inducing lepton asymmetry. We further assume that $N_{i}$ are somewhat heavier than the elementary fermion $n$. The part of the CHM3 Lagrangian, which describes the interactions of $n=N_{0}$ and $N_{i}$ with the SM leptons, $\tilde{L}_{2i}$ and Higgs doublet $H$, can be presented as a sum
@@
where $i,j=1,2,3$ and $x=0,1,2,3$. Because $\ell_{i}$ and $N_{0}$ are mainly elementary fields $|g_{i0}|\ll|g_{ij}|\lesssim|h_{j0}|\ll|h_{ij}|$. At the same time the Yukawa couplings $h_{ij}$ are not suppressed and therefore $|h_{ij}|\gtrsim 0.1$.
@@
It is convenient to introduce $Y_{\Delta B}$ which is the baryon asymmetry relative to the entropy density. The observed value of $Y_{\Delta B}$ is given by (see, for example, [42])
@@
In order to compute $Y_{\Delta B}$ the system of Boltzmann equations, that permit to calculate the evolution of the number density of $n$ and (BL)(B-L) number asymmetry generated by the decays of $n$, must be solved. The corresponding differential equations are somewhat similar to the Boltzmann equations describing thermal leptogenesis in the SM which are specified in Refs. [42, 43]. In the first approximation the induced baryon asymmetry can be estimated using an approximate formula [42, 43]
$Y_{\Delta B}\sim 10^{-3}\,\varepsilon_{0}\,\eta_{0}\,.$ (23)
In Eq. (23) $\eta_{0}$ is an efficiency factor that varies from 0 to 1. Without washout effects the decays of the Majorana fermion $n$ would result in $\eta_{0}=1$. The washout processes reduce the generated baryon asymmetry by the factor $\eta_{0}$.
For temperatures $T\lesssim 10\,\mbox{TeV}$, that correspond to the times $t\gtrsim 10^{-14}$ sec., the $(B+L)$ violating rate mediated by sphalerons is getting suppressed. Here we assume that the Majorana fermions $n$ decay before the sphaleron processes shut off, i.e. $H(T=M_{1})\lesssim{\Gamma}$, where $H$ is the Hubble expansion rate and $\Gamma$ is the total decay rate of $n$. In this so–called strong washout scenario the efficiency factor $\eta_{0}$ can be estimated as [43]
$\eta_{0}\simeq\left(\dfrac{H(T=M_{1})}{2\Gamma}\right)^{1.16}\,$$\qquad\qquad$
@@
where $g_{*}=n_{b}+\dfrac{7}{8}\,n_{f}$ is the number of relativistic degrees of freedom in the thermal bath. In the scenario under consideration $g_{*}=128.75$.
The requirement $H(T=M_{1})\lesssim{\Gamma}$ implies that $|h_{10}|\gtrsim 10^{-6}$. According to Eqs. (24), for such values of $|h_{10}|$ the efficiency factor $\eta_{0}$ diminishes with increasing $|h_{10}|$. The corresponding dependence is shown in Fig. 1 . In Eq. (23) both the decay asymmetry $\varepsilon_{0}$ and the efficiency factor are smaller than unity. Thus the appropriate baryon asymmetry can be generated only if $\eta_{0}\gtrsim 10^{-7}$. The results of the numerical analysis presented in Fig. 1 demonstrate that for $M_{n}\simeq 10\,\mbox{TeV}$ this condition can be fulfilled when $|h_{10}|\ll 0.001$. Otherwise the absolute value of $Y_{\Delta B}$ tends to be too small. The relatively small $|h_{10}|\lesssim 10^{-5}$ needed for successful leptogenesis might be an indication that the Lagrangian of the CHM3 possesses an approximate $Z^{\prime}_{2}$ symmetry, under which all multiplets except $n$ are even while $n$ is $Z^{\prime}_{2}$ odd.
[…]
Figure 1: Logarithm (base 10) of the absolute values of the efficiency factor $\eta_{0}$ (Left) and decay asymmetry $\varepsilon_{0}$ (Right) for $h_{20}=h_{30}=0$. The absolute value of $\eta_{0}$ is given as a function of logarithm (base 10) of $|h_{10}|$ for $M_{n}=10\,\mbox{TeV}$ (solid line) and $M_{n}=15\,\mbox{TeV}$ (dashed line). The absolute value of $\varepsilon_{0}$ is presented as a function of logarithm (base 10) of $\Delta\varphi_{1}=\Delta\varphi_{2}=\Delta\varphi_{3}=\Delta\varphi$ for $M_{n}=10\,\mbox{TeV}$, $M_{1}=12\,\mbox{TeV}$, $M_{2}=14\,\mbox{TeV}$, $M_{3}=16\,\mbox{TeV}$, $|h_{11}|=|h_{12}|=|h_{13}|=|h|$, $|h|=0.1$ (solid line) and $|h|=0.3$ (dashed line).
Since in the scenario under consideration the efficiency factor may not be negligibly small, i.e. $\eta_{0}\sim 0.01$, the absolute value of $Y_{\Delta B}$ is determined by the decay asymmetry $\varepsilon_{0}$. From Eq. (21) one can see that $\varepsilon_{0}$ is set by $|h_{1j}|$ and CP violating phases $\Delta\varphi_{j}$. It also depends on the ratio of masses of Majorana fermions $M_{j}/M_{n}$. Because the Yukawa couplings of $N_{j}$ to $L_{2i}$ and the Higgs doublet are not suppressed, all $|h_{ij}|$ should be relatively large, i.e. $|h_{ij}|\gtrsim 0.1$. Here we fix $M_{n}=10\,\mbox{TeV}$, $M_{1}=12\,\mbox{TeV}$, $M_{2}=14\,\mbox{TeV}$, $M_{3}=16\,\mbox{TeV}$, $|h_{11}|=|h_{12}|=|h_{13}|=h$ and $\Delta\varphi_{1}=\Delta\varphi_{2}=\Delta\varphi_{3}=\Delta\varphi$. In Fig. 1 the dependence of $|\varepsilon_{0}|$ on $\Delta\varphi$ for different values of $h$ is shown. The absolute value of the decay asymmetry grows monotonically with increasing of $|h|$. The results presented in Fig. 1 indicate that for $h\gtrsim 0.1$ and $\eta_{0}\sim 0.01$ the observed baryon asymmetry can be obtained for small CP violating phases, i.e. $\Delta\varphi\ll 0.01$.